# Learning functions on symmetric matrices and point clouds via lightweight invariant features

### v2

@@ 6 Extension of DeepSets with invariant features
where $\sigma:\mathbb{R}^{s}\to\mathbb{R}^{k}$ and $\phi:\mathbb{R}^{d}\to\mathbb{R}^{s}$ are typically (trained) standard machine learning models such as multi-layer perceptrons (denoted as MLP below), and $I$ is an index set for the (finite) input multiset. The function $\phi$ is often called the encoder, $\rho$ is the decoder, and $\mathbb{R}^{s}$ is the latent space. This machine learning model was introduced in [ZKR+17] and widely applied. Its mathematical properties, such as universality, have recently been studied [WFE+22, TW24, AGA+24].
In particular, it follows from [TW24] that there exists a family of specific “universal” encoders $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{s}$ with $s=2nd$ (where $n$ is an upper bound on the number of elements in the input set, and $d$ is the dimension of each of those elements) such that in the resulting DeepSets, the map $\mathbb{R}^{d\times n}\rightarrow\mathbb{R}^{s}$ to the latent space given by $V\mapsto\sum_{v\in\operatorname{col}(V)}\phi(v)$ is a permutation separator in the sense of Section 5. It follows that these DeepSets are universal approximators of $\mathfrak{S}_{n}$-invariant continuous functions in compact sets of $\mathbb{R}^{d\times n}$. This is interpreted as implying that DeepSets with at worst $O(n)$ latent dimension are universally expressive of $\mathfrak{S}_{n}$-invariant functions. That said, from a practical standpoint, when learning a single function with low-dimensional output, it is overly constraining to commit at the outset to one of the “universal” encoders $\phi$ with its latent space dimension of $2nd$. Instead, the models used in practice learn both $\sigma$ and $\phi$ from data, and may set $s\ll 2nd$, especially when the training data set is small.
We propose to use DeepSets in combination with the results of Sections 3, 4, and 5. For symmetric matrix data, rather than computing the $f$’s from the input $X$ by sorting, we use the sets $\{X_{ii}\}$ and $\{X_{ij}\}$ as inputs to a DeepSet model, dubbed DeepSet for Conjugation Invariance (DS-CI):
@@
In the application to point clouds, we implement the permutation separators $D$ as DeepSets. In particular, we propose $\mathrm{O}(d)$-Invariant DeepSet (OI-DS) as follows:
@@ 7.1 Molecular property prediction on QM7b
Table 1: We report the mean absolute error (MAE) on the test set over 10 random data splits (80/10/10 for train/validation/test sets). The results for the Kernel Ridge Regression (KRR) and Deep Tensor Neural Network (DTNN) with the same data split ratio are taken from [WRF+18, Table 10].
@@ 7.2 Point cloud distance prediction on ModelNet10
We investigate both the DS-CI model (5) with all $f$s (middle subplots) and the OI-DS model (6), using $k$-means centroids as the equivariant identifiers (right subplots) with $k=d=3$. As shown in Figure 2 below, both models successfully predict the GW-based distances, and the performance is robust to replacing $O(n^{2})$ invariant features in DS-CI with $O(n)$ features (for fixed $d$) in OI-DS.
Table 2: Rank correlation of the target and predicted distances
Rank Corr. $\uparrow$ DS-CI OI-DS
Training set $0.88$ $0.86$
Test set $0.80$ $0.75$
@@ 8 Discussion
In this work, we provide $O(n^{2})$ functions on $n\times n$ symmetric matrices that are invariant with respect to the action of permutations by conjugation, and separate orbits outside a closed measure zero subset of the space of symmetric $n\times n$ matrices (the “bad set”). Thus they can be used to universally approximate invariant functions outside the same bad set. We also show that when restricted to Gram matrices of point clouds in $\mathbb{R}^{d}$, these functions again separate orbits (this time for the natural action of $\mathrm{O}(d)\times\mathfrak{S}_{n}$) outside a “bad set” which is closed and measure zero inside the space of point clouds. Again, it follows that these functions can be used to approximate invariant functions on point clouds.
To improve efficiency for $d$ fixed and $n$ large, we show how to use $O(n)$ invariant features to approximate any continuous invariant functions on $d$-dimensional point clouds (again outside a “bad set” which is closed and measure zero). We demonstrate the feasibility of our approach via numerical experiments on molecular property regression and point cloud distance prediction.
@@ Acknowledgements
We thank Alexandra Pevzner (UMN) for research discussions at the early stages of this project, and for feedback on the proof of Proposition 4.2, including the main idea for the $d=1$ case. We thank David W. Hogg (NYU), Kate Storey-Fisher (DIPC), Gaby Contardo (SISSA), Francisco Villascusa (Simons Foundation), and Kaze Wong (Flatiron Institute) for giving us context about point clouds in cosmology that inspired this research; some of these conversations took place at the Coworking Retreat on Equivariant Machine Learning held at Johns Hopkins University in March 2023. We also thank Puoya Tabaghi (UCSD) for conversations about [TW24]; Facundo Mémoli (OSU) for conversations about the Gromov-Wasserstein lower bounds; Daniel Packer (OSU) for implementing the Gromov-Wasserstein lower bound for OTT-JAX that we use in our experiments; Tim Carson (Google) for conversations about $k$-means clustering; and Erik Thiede (Cornell) for suggestions on the molecule property regression experiments. SV and BBS are partially supported by ONR N00014-22-1-2126 and NSF CCF 2212457. SV is also funded by the NSF–Simons Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning (MoDL) (NSF DMS 2031985), and NSF CAREER 2339682.
@@ Appendix A Appendix
We give the proofs of Proposition 3.4 and Proposition 4.2. For Proposition 3.4, recall that the proposition states that the $f^{d}$’s, the $f^{o}$’s, and $f^{\star}$, defined in Section 3, separate orbits for the action of the symmetric group $\mathfrak{S}_{n}$ on the space $\mathcal{S}(n)$ of symmetric matrices, away from a closed, $\mathfrak{S}_{n}$-stable, measure zero subset $B$ of “bad” matrices. The argument makes use of the elementary symmetric polynomials $e^{d}_{1}(X),\dots,e^{d}_{n}(X)$ in $X_{11},\dots,X_{nn}$, defined by the following formal identity:
@@
It follows that the $f^{d}$’s, the $f^{o}$’s, and $f^{\star}$ (taken as a whole) distinguish a pair of orbits if and only if the $e^{d}$’s, the $e^{o}$’s, and $f^{\star}$ (taken as a whole) distinguish the same pair of orbits. Thus, it suffices to prove the proposition with $e^{d}$’s and $e^{o}$’s in the place of $f^{d}$’s and $f^{o}$’s respectively, and this is what we will do.
@@
The proposition now follows immediately from [BBSK+23, Corollary 3.31] (applied with $U$ as the linear span of the $e^{d}$’s, the $e^{o}$’s, and $f^{\star}$). The conclusion is that the map
@@
We turn to Proposition 4.2. The idea is essentially the same as for Proposition 3.4; however, we need a little bit of additional work to handle the fact that the Gram matrices of point clouds of dimension $d<n$ belong to a small proper subvariety of $\mathcal{S}(n)$, in general much smaller than the bad set $B$ of Proposition 3.4.
Proof of Proposition 4.2.
@@
The $\tilde{e}$’s contain the same information, and thus have the same power to separate orbits, as the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s. Thus, if we can show Proposition 4.2 holds with the $\tilde{e}$’s in the place of the $\tilde{f}^{o}$’s and $\tilde{f}^{d}$’s, then it also holds as written.
To obtain a version of Proposition 4.2 with the $\tilde{e}$’s in place of the $\tilde{f}^{o}$’s and $\tilde{f}^{d}$’s, our goal will be to show that $\tilde{f}^{\star}$ and the $\tilde{e}$’s generate the field of rational functions on $\mathbb{R}^{d\times n}$ that are invariant under the $\mathrm{O}(d)\times\mathfrak{S}_{n}$ action. The desired statement—i.e., the existence of a closed, measure-zero set $\tilde{B}\subset\mathbb{R}^{d\times n}$ on the complement of which any two $\mathrm{O}(d)\times\mathfrak{S}_{n}$-orbits are separated by some $\tilde{f}$—then follows by [BBSK+23, Corollary 3.31] (or by Rosenlicht’s theorem [Ros56, Theorem 2], or by the proof sketch in Remark A.3, which goes through unchanged in this setting).
@@
The action of $\mathfrak{S}_{n}$ on $L$ in the previous paragraph is by simultaneous row and column permutations of the matrix $\overline{X}:=(\overline{X}_{ij})$. In the proof of Proposition 3.4 (in particular, in Lemma A.1), we used the symbol $G$ to refer to $\mathfrak{S}_{n}$ in the analogous action on $R$, to distinguish it as a particular subgroup of the permutation group $\Gamma\cong\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ which acts independently on the $X_{ii}$’s and $X_{ij}$’s. We retain that notation here; in this language, what we must do is to show that $\overline{f}^{\star}$ and the $\overline{e}$’s generate $L^{G}$ as a field over $\mathbb{R}$.
@@
We thank Alexandra Pevzner for the main idea for handling the $d=1$ case in the above.

### v3

@@ 7 Extension of DeepSets with invariant features
where $\sigma:\mathbb{R}^{s}\to\mathbb{R}^{k}$ and $\phi:\mathbb{R}^{d}\to\mathbb{R}^{s}$ are typically (trained) standard machine learning models such as multi-layer perceptrons (denoted as MLP below), and $I$ is an index set for the (finite) input multiset. The function $\phi$ is often called the encoder, $\sigma$ is the decoder, and $\mathbb{R}^{s}$ is the latent space. This machine learning model was introduced in [ZKR+17] and widely applied. Its mathematical properties, such as universality, have recently been studied [WFE+22, TW24, AGA+24].
In particular, it follows from [TW24] that there exists a family of specific “universal” encoders $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{s}$ with $s=2nd$ (where $n$ is an upper bound on the number of elements in the input set, and $d$ is the dimension of each of those elements) such that in the resulting DeepSets, the map $\mathbb{R}^{d\times n}\rightarrow\mathbb{R}^{s}$ to the latent space given by $V\mapsto\sum_{v\in\operatorname{cols}(V)}\phi(v)$ is a permutation-orbit separator in the sense of Section 6. It follows that these DeepSets are universal approximators of $\mathfrak{S}_{n}$-invariant continuous functions in compact sets of $\mathbb{R}^{d\times n}$. This is interpreted as implying that DeepSets with at worst $O(n)$ latent dimension are universally expressive of $\mathfrak{S}_{n}$-invariant functions. That said, from a practical standpoint, when learning a single function with low-dimensional output, it is overly constraining to commit at the outset to one of the “universal” encoders $\phi$ with its latent space dimension of $2nd$. Instead, the models used in practice learn both $\sigma$ and $\phi$ from data, and may set $s\ll 2nd$, especially when the training data set is small.
We propose to use DeepSets in combination with the results of Sections 4, 5, and 6. For symmetric matrix data, rather than computing the $f$’s from the input $X$ by sorting, we use the sets $\{X_{ii}\}$ and $\{X_{ij}\}$ as inputs to a DeepSet model, dubbed DeepSet for Conjugation Invariance (DS-CI):
@@
In the application to point clouds, we implement the permutation-orbit separators $D$ as DeepSets. In particular, we propose $\mathrm{O}(d)$-Invariant DeepSet (OI-DS) as follows:
@@ 8.1 Molecular property prediction on QM7b
Table 1: We report the mean absolute error (MAE) on the test set over 10 random data splits (80/10/10 for train/validation/test sets). The results for the Kernel Ridge Regression (KRR) and Deep Tensor Neural Network (DTNN) with the same data split ratio are taken from [WRF+18, Table 10]. Bold numbers indicate the best results.
@@ 8.2 Point cloud distance prediction on ModelNet10
We investigate both the DS-CI model (6) with all $f$s (middle subplots) and the OI-DS model (7), using $k$-means centroids as the equivariant identifiers (right subplots) with $k=d=3$. As shown in Figure 2 and Table 2, both models predict the GW-based distances reasonably well, evidenced by a high rank correlation and low mean squared error. While the mean relative error $\frac{1}{n}\sum_{i=1}^{n}|\text{GW}_{i}-\widehat{\text{GW}}_{i}|/\text{GW}_{i}$ is relatively high, it is inflated due to many target distances $\text{GW}_{i}$ being close to zero. The performance degrades mildly when replacing $O(n^{2})$ invariant features in DS-CI with $O(n)$ features (for fixed $d$) in OI-DS.
Table 2: Rank correlation, mean squared error (MSE), and relative error of the target and predicted distances.
DS-CI OI-DS
Rank Corr. $\uparrow$ MSE $\downarrow$ Rel. Error $\downarrow$ Rank Corr. $\uparrow$ MSE $\downarrow$ Rel. Error $\downarrow$
Training set $0.88$ $0.01$ $0.53$ $0.75$ $0.04$ $0.88$
Test set $0.86$ $0.02$ $0.60$ $0.73$ $0.03$ $0.63$
@@ 9 Discussion
In this work, we provide a general method to extend a set of generically separating invariants of a Lie group acting on a topological measure space by isometries, to a set of generically separating invariants of a subgroup of finite index. The construction is inspired by ideas from Galois theory, and it applies to classes of functions commonly used in machine learning.
Using this construction, we provide $O(n^{2})$ functions on $n\times n$ symmetric matrices that are invariant with respect to the action of permutations by conjugation, and separate orbits outside a closed measure zero subset of the space of symmetric $n\times n$ matrices (the “bad set”). Thus they can be used to universally approximate continuous invariant functions in compact sets outside the same bad set. We also show that when restricted to Gram matrices of point clouds in $\mathbb{R}^{d}$, these functions again separate orbits (this time for the natural action of $\mathrm{O}(d)\times\mathfrak{S}_{n}$) outside a “bad set” which is closed and measure zero inside the space of point clouds. Again, it follows that these functions can be used to approximate invariant continuous functions on point clouds.
To improve efficiency for $d$ fixed and $n$ large, we show how to use $O(n)$ invariant features to approximate any continuous invariant functions on $d$-dimensional point clouds (again in compact sets outside a “bad set” which is closed and measure zero). We demonstrate the feasibility of our approach via numerical experiments on molecular property regression and point cloud distance prediction.
@@ Acknowledgements
We thank Alexandra Pevzner (UMN) for research discussions at the early stages of this project, and for feedback on the proof of Proposition 5.2, including the main idea for the $d=1$ case. We thank David W. Hogg (NYU), Kate Storey-Fisher (DIPC), Gaby Contardo (SISSA), Francisco Villascusa (Simons Foundation), and Kaze Wong (Flatiron Institute) for giving us context about point clouds in cosmology that inspired this research; some of these conversations took place at the Coworking Retreat on Equivariant Machine Learning held at Johns Hopkins University in March 2023. We also thank Puoya Tabaghi (UCSD) for conversations about [TW24]; Facundo Mémoli (OSU) for conversations about the Gromov-Wasserstein lower bounds; Daniel Packer (OSU) for implementing the Gromov-Wasserstein lower bound for OTT-JAX that we use in our experiments; Tim Carson (Google) for conversations about $k$-means clustering; and Erik Thiede (Cornell) for suggestions on the molecule property regression experiments. SV and BBS were partially supported by ONR N00014-22-1-2126, NSF CCF 2212457, and NSF CAREER 2339682. SV is also funded by the NSF–Simons Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning (MoDL) (NSF DMS 2031985), and NSF 2430292.
@@ Appendix A Appendix
We give the proof of Theorem 5.2, and the original (Galois-theoretic) proof of Theorem 4.4. For Theorem 4.4, recall that the theorem states that the $f^{d}$’s, the $f^{o}$’s, and $f^{\star}$, defined in Section 4, separate orbits for the action of the symmetric group $\mathfrak{S}_{n}$ on the space $\mathcal{S}(n)$ of symmetric matrices, away from a closed, $\mathfrak{S}_{n}$-stable, measure zero subset $B$ of “bad” matrices. The argument makes use of the elementary symmetric polynomials $e^{d}_{1}(X),\dots,e^{d}_{n}(X)$ in $X_{11},\dots,X_{nn}$, defined by the following formal identity:
@@
It follows that the $f^{d}$’s, the $f^{o}$’s, and $f^{\star}$ (taken as a whole) distinguish a pair of orbits if and only if the $e^{d}$’s, the $e^{o}$’s, and $f^{\star}$ (taken as a whole) distinguish the same pair of orbits. Thus, it suffices to prove the theorem with $e^{d}$’s and $e^{o}$’s in the place of $f^{d}$’s and $f^{o}$’s respectively, and this is what we will do.
@@
The theorem now follows immediately from [BBSK+23, Corollary 3.31] (applied with $U$ as the linear span of the $e^{d}$’s, the $e^{o}$’s, and $f^{\star}$). The conclusion is that the map
@@
We turn to Theorem 5.2. The idea is essentially the same as for Theorem 4.4; however, we need a little bit of additional work to handle the fact that the Gram matrices of point clouds of dimension $d<n$ belong to a small proper subvariety of $\mathcal{S}(n)$, in general much smaller than the bad set $B$ of Theorem 4.4.
Proof of Theorem 5.2.
@@
For example,
$\tilde{f}^{\star}=\sum_{i\neq j}(v_{i}^{\top}v_{i})(v_{i}^{\top}v_{j}),$
and
$\tilde{e}^{d}_{3}=\sum_{i<j<k}(v_{i}^{\top}v_{i})(v_{j}^{\top}v_{j})(v_{k}^{\top}v_{k}).$
The $\tilde{e}$’s contain the same information, and thus have the same power to separate orbits, as the $\tilde{f}^{d}$’s and $\tilde{f}^{o}$’s. Thus, if we can show Theorem 5.2 holds with the $\tilde{e}$’s in the place of the $\tilde{f}^{o}$’s and $\tilde{f}^{d}$’s, then it also holds as written.
To obtain a version of Theorem 5.2 with the $\tilde{e}$’s in place of the $\tilde{f}^{o}$’s and $\tilde{f}^{d}$’s, our goal will be to show that $\tilde{f}^{\star}$ and the $\tilde{e}$’s generate the field of rational functions on $\mathbb{R}^{d\times n}$ that are invariant under the $\mathrm{O}(d)\times\mathfrak{S}_{n}$ action. The desired statement—i.e., the existence of a closed, measure-zero set $\tilde{B}\subset\mathbb{R}^{d\times n}$ on the complement of which any two $\mathrm{O}(d)\times\mathfrak{S}_{n}$-orbits are separated by some $\tilde{f}$—then follows by [BBSK+23, Corollary 3.31] (or by Rosenlicht’s theorem [Ros56, Theorem 2], or by the proof sketch in Remark A.3, which goes through unchanged in this setting).
@@
The action of $\mathfrak{S}_{n}$ on $L$ in the previous paragraph is by simultaneous row and column permutations of the matrix $\overline{X}:=(\overline{X}_{ij})$. In the proof of Theorem 4.4 (in particular, in Lemma A.1), we used the symbol $G$ to refer to $\mathfrak{S}_{n}$ in the analogous action on $R$, to distinguish it as a particular subgroup of the permutation group $\Gamma\cong\mathfrak{S}_{n}\times\mathfrak{S}_{n(n-1)/2}$ which acts independently on the $X_{ii}$’s and $X_{ij}$’s. We retain that notation here; in this language, what we must do is to show that $\overline{f}^{\star}$ and the $\overline{e}$’s generate $L^{G}$ as a field over $\mathbb{R}$.
@@
We isolate the invariant-theoretic content of this proof:
Corollary A.4.
The $\tilde{e}$’s and $\tilde{f}^{\star}$ generate the field
$\mathbb{R}\left(\mathbb{R}^{d\times n}\right)^{\mathrm{O}(d)\times\mathfrak{S}_{n}}$
of rational functions on $\mathbb{R}^{d\times n}$ invariant under the left action by $\mathrm{O}(d)$ and the right action by $\mathfrak{S}_{n}$.∎
We thank Alexandra Pevzner for the main idea for handling the $d=1$ case in this proof .